ax²+bx+c 01 x= -htb² - 4ac Ex3 - 2x² +3 x =4-15 X X - 2x²-x =-15 +15 +15 - 2x2-x+15= 0 A = -2 D- C= 15 algebra 2 Solving the quadratic equation using the quadratic formula
Answer #1

Solution => – 2012 – X+15= 0 A=-2 B = -1 c=15 Tv X=-bI/6²-4ac za This is the formula for finding the doors of a quadoatic equcase 1-11=-105 E=4 -- - Therefore, the root of the quadralit equation .-200² _ X+15=0 are 9 and 5/2 X=-3, 5/2